OpenStudy (anonymous):

There is one critical point for z = x^5y + xy^5 + xy. Find it.

OpenStudy (anonymous):

can someone just explain how to get started??

OpenStudy (anonymous):

take partial dz/dx= 5x^4 y + y^5 +y dz/dy= x^5 +x 5y^4+x

OpenStudy (anonymous):

set partial=0

OpenStudy (anonymous):

ok and that is the critical point? seemed more confusing in the text. thanks!

OpenStudy (anonymous):

I got x=0 y=0

OpenStudy (anonymous):

x=0 y=0 makes both partial zero

OpenStudy (anonymous):

would the same method by applied here as well?? Find the critical point(s) of f(x, y) = 1 + (x − y)^2 + (x − 1)^2

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

ok thank you!!

OpenStudy (anonymous):

It is basically just like critical value in calc I but we have both x and y so it is called crictle points

OpenStudy (anonymous):

oh ok!